OpenStudy (anonymous):

Please help:) what is 6√150x simplified?

OpenStudy (accessdenied):

Well, if you break up 150 into a product of a perfect square, we get \( 6 \sqrt{150x} = 6 \sqrt{6 \times \mathbf{25} x} \) From there, you should be able to pull out a 5 from the property \(\sqrt{ab} = \sqrt{a}\sqrt{b}\).

OpenStudy (anonymous):

So it'd be 30x √ 6?

OpenStudy (anonymous):

I get confused with the variable

OpenStudy (accessdenied):

You cannot remove the \(x\) from the square root because it is only multiplied once. If it were something like \(x^2\), we could bring out an \(|x|\) because \(\sqrt{x^2} = |x|\) (the absolute values make sure that we respect that x can be positive or negative and still be \(x^2\))

OpenStudy (anonymous):

That makes sense! So then it would be \[30\sqrt{6x}\]

OpenStudy (accessdenied):

Yes. :)

OpenStudy (anonymous):

Thanks, I keep learning things every five seconds on open study:D